# QUIDS: Query Intent Generation via Dual Space Modeling

### v2

@@
Abstract.
Query understanding is a crucial component of Information Retrieval (IR), aimed at identifying the underlying search intent of textual queries. However, most existing approaches oversimplify this task into query classification or clustering, which fails to fully capture the nuanced intent behind the query. In this paper, we address the task of query intent generation: to automatically generate detailed and precise intent descriptions for search queries using relevant and irrelevant documents given a query. These intent descriptions can help users understand why the search engine considered the top-ranked documents relevant, and provide more transparency to the retrieval process. We propose a dual-space model that uses semantic relevance and irrelevance information in the returned documents to explain the understanding of the query intent. Specifically, in the encoding process, we project, separate, and distinguish relevant and irrelevant documents in the representation space. Then, we introduce a semantic decoupling model in the novel disentangling space, where the semantics of irrelevant information are removed from the relevant space, ensuring that only the essential and relevant intent is captured. This process refines the understanding of the query and provides more accurate explanations for the search results. Experiments on benchmark data demonstrate that our methods produce high-quality query intent descriptions, outperforming existing methods for this task, as well as state-of-the-art query-based summarization methods. A token-level visualization of attention scores reveals that our model effectively reduces the focus on irrelevant intent topics. Our findings open up promising research and application directions for query intent generation, particularly in exploratory search.
Information Retrieval, Query intent generation, Contrastive learning, Transformers
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: ; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Users and interactive retrievalccs: Information systems Query intentccs: Computing methodologies Natural language generation
1. Introduction
Query understanding is a focal point in the Information Retrieval (IR) community, aimed at determining the underlying search intent of a textual query input. This facet of IR is crucial for several reasons. Firstly, it enables search engines to distinguish and filter out unwanted interpretations of the current search query, ensuring relevance and accuracy in results. This is important because queries tend to be short and underspecified and are therefore often ambiguous (Guo and Lan, 2020). Secondly, users benefit when search engines provide feedback, such as query expansions or more descriptive query intents (Chen et al., 2021). This enhances the search experience by enabling users to refine and adjust their queries based on insights derived from the provided query understanding.
@@ 1. Introduction
Figure 1. An application of the query intent generation task as a form of explanation to the user in exploratory search. The top-ranked documents are used as pseudo-relevant documents to generate the intent. Information relevant to forming the intent is highlighted in boldface.
Existing methods for query understanding can be broadly categorized into three groups. Query classification methods (Broder, 2002; Verberne et al., 2013) aim to categorize queries into distinct classes based on intent taxonomies. However, many of these methods tend to abstract intents into broader categories, lacking granularity in their analysis. Second, query clustering methods (Wen et al., 2002; Hong et al., 2016) aim to group query intents into clusters, yet encounter challenges in precisely interpreting individual queries within these clusters. The third approach, which is the most effective way to capture query intent, is query intent description generation, where the intent is described in textual format. However, this task has been largely overlooked in existing work, with the most recent contribution by (Zhang et al., 2020), which still relies on RNN-based methods and does not leverage the advancements in transformer models.
Figure 1 illustrates an example of query intent generation for the query ‘Freon-12’, which has an exploratory intent – aimed at exploring topics without a clearly defined goal  (Palagi et al., 2017). Since the true intent behind the query is unknown to the search engine, it relies on the retrieved documents to infer it. Thus, it treats the top-ranked documents as pseudo-relevant and the low-ranked documents as irrelevant. The intent description generated from both sets of documents is highlighted with a green background, while the information directly contributing to intent formation is marked in boldface. From the relevant documents, key terms like ‘Freon-12’, ‘refrigerant’, and ‘phased out’ are captured and emphasized in the intent description. In contrast, topics such as ‘costs’ and ‘prices’, which appear in both relevant and irrelevant documents, are deprioritized and excluded from the final intent description. The resulting intent generation illustrates how the search engine interprets the user’s query based on the retrieved documents. Furthermore, this approach allows users to reflect on their interests more clearly, potentially guiding them to reformulate their query more precisely and achieve more targeted search results. This iterative process of query refinement helps bridge the gap between exploratory search and more specific information needs, enhancing the overall search experience.
In scenarios where the ground truth intent is available, a related task to train models for generating query intent is query-focused summarization (QFS). This task is to generate summaries tailored to specific user queries or information needs, drawing from pseudo-relevant (high-ranked) documents. However, the focus in QFS is on summarizing content rather than explicitly identifying or reproducing query intents. Typically, QFS does not involve irrelevant information, which means the resulting summaries may include unfiltered or irrelevant details that are not truly reflective of the query’s intent. On the other hand, previous work (Zhang et al., 2020) on the Q2ID task incorporates irrelevant documents for contrast, but it does not establish a direct connection between the query and irrelevant documents. Our approach explicitly models both relevant and irrelevant intent spaces in a query-aware manner. Specifically, we construct the final intent space by subtracting the irrelevant intent space from the relevant one. This subtraction ensures that the final intent description is closely aligned with the query while effectively filtering out irrelevant information.
We propose a novel pipeline, implementing contrastive learning through dual space modeling in transformer-based (Vaswani et al., 2017) models, particularly T5 (Raffel et al., 2020a) and BART (Lewis et al., 2020). Specifically, we use a dual-encoder architecture and perform contrastive learning in the representation space using encoder embeddings of both relevant and irrelevant documents. During the decoding phase, we adapt the transformer decoder to conduct contrastive learning using decoder hidden states in the newly proposed disentangling space, where we model query-aware relevant and irrelevant intent spaces. This dual-space contrastive learning approach ensures that the model learns to differentiate between relevant and irrelevant documents, while also guiding it to produce high-quality intent descriptions with attention to the learned relevant and irrelevant intent space. Moreover, we introduce a method to augment hard negative documents driven by intent, to expand the irrelevant information representation space during training.
We evaluate our model using the benchmark dataset Q2ID, sourced from TREC and SemEval, which includes two types of query intent. Our results demonstrate that our model outperforms state-of-the-art baselines in terms of ROUGE metrics and BERTScore. Additionally, we perform human evaluations and LLM-based evaluations using tailored metrics, and provide a comprehensive analysis of our approach. To further illustrate the effect of irrelevant documents, we conduct a case study in which we highlight the information used to generate a semantically rich intent token based on the decoder’s cross-attention weights at the token level for both relevant and irrelevant documents. In summary, our contributions are as follows:
  • We introduce contrastive learning in both the representation space and the disentangling space of transformer models, effectively capturing contrasting information from relevant and irrelevant documents.
  • Our model generates high-quality intent descriptions, with performance significantly enhanced by incorporating hard negative data augmentation during training.
  • We perform a thorough evaluation of our model, providing us with insights into the model’s strengths and weaknesses, as well as its potential application scenarios.
2. Related Work
2.1. Query Understanding
The most closely related tasks in query understanding include query classification, query clustering, and query expansion. Query classification (Guo and Lan, 2020; Li et al., 2023) aims to classify queries into one or more pre-defined target taxonomies, based on factors like intent (Broder, 2002; Rose and Levinson, 2004), topic (Li et al., 2005; Beitzel et al., 2005), information type (Pasca and Harabagiu, 2001) or other classification dimensions. However, due to the typically short and ambiguous nature of search queries, as well as their evolving meanings, these taxonomies may lack sufficient granularity and become outdated over time (Guo and Lan, 2020). This can lead to the loss of detailed information in the classification process. Query clustering attempts to discover search topics or intents by mining clusters of queries. The similarity between queries is typically measured using data from user click-through logs (Wen et al., 2002; Rani and Babu, 2019; Lu et al., 2024) or top-ranked documents (Hong et al., 2016). In a recent application, query clustering has been employed in search engines to display real-time trending user queries, capturing shifts in queries and news articles over consecutive timestamps (Lu et al., 2024). Earlier studies on question answering demonstrate that an effective way to expand a query is to extract answer patterns or select terms that could be possible answers as expansion terms. Recently, some generation-augmented retrieval methods (Mao et al., 2021; Chen et al., 2022) focus on exploiting the knowledge captured in pre-trained language models (PLMs) (Roberts et al., 2020; Brown et al., 2020) to generate the potential answer as expansion terms.
Inspired by benchmark datasets containing detailed descriptions provided by human annotators, Zhang et al. (2020) propose the Query-to-Intent-Description (Q2ID) task that aims to generate a natural language intent description based on both relevant and irrelevant documents of a given query. Unlike their method, we directly model a query-aware irrelevant intent space in transformer models, leading to a more precise intent description.
2.2. Query-Focused Summarization
Query-focused summarization (QFS) is a subtask within text summarization that aims to generate a summary of a given text based on a specific query. Traditionally, the predominant approaches are unsupervised extraction methods (Wan and Xiao, 2009; Feigenblat et al., 2017) that generate summaries by ranking and extracting textual segments according to their similarity to other segments and their relevance to the query. In the past years, numerous QFS datasets (Kulkarni et al., 2020; Fabbri et al., 2022; Zhong et al., 2021) have been introduced, primarily based on answer selection or summaries for specific queries, leading to the development of many recent QA-motivated methods (Su et al., 2020, 2021; Egonmwan et al., 2019). Other works focus on query modeling through approaches such as the learning evidence ranking model (Xu and Lapata, 2021), jointly optimizing a latent query model and a conditional language model (Xu and Lapata, 2022), or designing pipeline models for abstractive summarizations. For instance, (Xu and Lapata, 2020) introduce a coarse-to-fine pipeline consisting of three estimators (relevance, evidence, and centrality estimators) that gradually discard less relevant segments.
Effectively extracting key insights from long documents has become a critical challenge. Vig et al. (2022) evaluate a two-step extract-then-abstract approach, where a scoring model is trained to select relevant portions of the text based on a query, which are then used as input to an abstraction model to generate a final summary. Additionally, they propose end-to-end methods that handle longer input texts by employing sparse attention mechanisms. Additional approaches focus on the use of a question-driven pretraining objective (Pagnoni et al., 2023), contrastive learning based on the scored segments (Sotudeh and Goharian, 2023), and joint modeling of token and utterance based on query-utterance attention (Liu et al., 2023a). In this work, we use state-of-the-art QFS methods as baselines for our proposed model.
3. Methods
@@
Figure 2. Overview of our proposed pipeline. From left to right, we show (a) Intent-Driven Negative Augmentation method, (b) Contrastive decoder structure with dual cross-attention layers, (c) and (d) Contrastive learning via dual space modeling.
@@
In this section, we present our approach QUIDS that generates descriptive query intents via dual space modeling. Figure 2 shows an overview. We first introduce the pipeline framework in section 3.1, then the data augmentation method to select intent-driven hard negative samples in section 3.2, and finally the contrastive mechanism via dual space modeling in section 3.3.
3.1. Pipeline Framework
We define the contrastive generation task as follows. Given a dataset $\mathcal{D}=\{(q,\mathcal{R},\mathcal{I},y)_{j}\}$ with $L$ samples, where $j\in\{0,1,...,L\}$: $q$ is a query, =$\mathcal{R}=\{r_{1},r_{2},...,r_{|\mathcal{R}|}\}$ is a collection of relevant documents for the query, =$\mathcal{I}=\{i_{1},i_{2},...,i_{|\mathcal{I}|}\}$ is a collection of irrelevant documents and $y$ is the ground truth query intent (i.e. the expected output). The goal is to learn the distinctions between relevant and irrelevant inputs based on a query while generating an intent description that exclusively highlights the relevant aspects related to the query. To achieve this, our training pipeline consists of 2 steps: (1) Intent-Driven Negative Augmentation (IDNA): mining hard negative documents as irrelevant documents from the entire dataset $D$ based on the query, its relevant document collections, and the ground truth intent, i.e.,
$IDNA(q,\mathcal{R},y,\mathcal{D})=\mathcal{I}^{\prime}$
@@ 3.1. Pipeline Framework
where =$\mathcal{I}^{\prime}=\{i^{\prime}_{1},i^{\prime}_{2},...,i^{\prime}_{h}\}$ with $h$ the expected number of irrelevant documents. (2) Dual Space Modeling (DSM): contrastively generating a descriptive intent for the query by modeling query-aware relevant and irrelevant intent spaces, i.e.,
$DSM(q,\mathcal{R},i^{\prime})=\hat{y}$
3.2. Intent-Driven Negative Augmentation
Inspired by Liu et al. (2022) on choosing in-context sample strategies for in-context learning, we design a method to choose intent-aware hard negative samples based on semantic similarity. Specifically, we use a Sentence Transformer model (Reimers and Gurevych, 2019) to represent both positive and negative samples from the training data in a vector space. Then we choose negative samples close to the positive ones in this space. The negative sample augmentation makes the task more challenging for the model, hence improving its discriminative capabilities. As shown in Figure 2 (a), we augment hard negative samples using the following steps: For each query $q$ in the training set,
  1. (1)
    We rank its relevant documents $\mathcal{R}$ based on their cosine similarity to the concatenated embedding of the query and its ground truth intent $(q;y)$ where (;) denotes concatenation, resulting in $\mathcal{R^{*}}$.
  2. (2)
    We concatenate the ranked relevant documents $\mathcal{R^{*}}$ into a single document $R^{*}$ and obtain its embedding $h_{R^{*}}$.
  3. (3)
    The original irrelevant documents are initially ranked based on their cosine similarities to the newly concatenated relevant embedding $h_{R^{*}}$.
  4. (4)
    If the number of irrelevant documents is smaller than an expected size, the irrelevant document collection $\mathcal{I}^{\prime}$ is augmented by selecting hard negative samples from all other documents in the whole dataset $\mathcal{D}$ according to their cosine similarities to the new embedding $h_{R^{*}}$.
For step 1, 3 and 4, we use a pre-trained Sentence Transformers model (Reimers and Gurevych, 2019) trained on the MSMARCO Passage Ranking Dataset (Nguyen et al., 2016) to encode the documents. In practice, during step 4, we select documents whose cosine similarities exceed a predefined threshold 0.8 until the expected size is reached.
3.3. Dual Space Modeling
@@
We use a Transformer-based encoder-decoder architecture, with the BART-large model (Lewis et al., 2020) and the T5-large model (Raffel et al., 2020a)as the backbone, as illustrated in Figure 2 (b). To capture the relationship between a query and its relevant and irrelevant documents, we adopt a Siamese dual encoder architecture. Based on the encoder outputs, contrastive learning is performed in the representation space to differentiate between embeddings for relevant and irrelevant documents. Correspondingly, we design a contrastive decoder to model query-aware relevant and irrelevant intent spaces in a disentangled manner.
3.3.1. Representation space modeling
Naturally, we expect the encoder to model and distinguish the relevant documents and irrelevant documents concerning a query in a feature representation space. To this end, we implement a dual encoder architecture, with each encoder being a cross-encoder, so that each cross-encoder directly models the relevance of a relevant or irrelevant document given a query by jointly encoding the query-document pair.
Intuitively, relevant documents are assumed to share similar topics, reflecting their connection to the specific information need behind a query to varying degrees. Therefore, it is natural to encode relevant documents collectively by concatenating them according to the ranking outlined in Section 3.2, Step 1. In contrast, irrelevant documents can be irrelevant to a query in diverse ways, making it impractical to model a meaningful and comprehensive irrelevant feature representation space. Instead, we focus on a single irrelevant document $i^{\prime}$ at each training step, using a hard negative sample from the augmented irrelevant document collection $\mathcal{I^{\prime}}$. We aim to model a feature representation space for relevant documents, reflecting their shared topics, while ensuring the irrelevant feature representation space is distant from the relevant one due to the dissimilarity of their topics. To achieve this, we project both of the encoder last hidden states into a feature representation space via a linear layer. The feature representation for each relevant document is obtained by applying average pooling over the token representations corresponding to the document. We then pull the relevant feature representations closer together while pushing the irrelevant representation further away from each relevant one (Figure 2 (c)). The optimization objective of the relevant feature representation space is to minimize:
(1) $\mathcal{L}_{rel}=\sum^{k}_{m=1}\sum^{k}_{n=m+1}d(e_{m},e_{n})$

### v3

# QUIDS: Query Intent Description for Exploratory Search via Dual Space Modeling

@@
Abstract
In exploratory search, users often submit vague queries to investigate unfamiliar topics, but receive limited feedback about how the search engine understood their input. This leads to a self-reinforcing cycle of mismatched results and trial-and-error reformulation. To address this, we study the task of generating user-facing natural language query intent descriptions that surface what the system likely inferred the query to mean, based on post-retrieval evidence. We propose QUIDS, a method that leverages dual-space contrastive learning to isolate intent-relevant information while suppressing irrelevant content. QUIDS combines a dual-encoder representation space with a disentangling decoder that works together to produce concise and accurate intent descriptions. Enhanced by intent-driven hard negative sampling, the model significantly outperforms state-of-the-art baselines across ROUGE, BERTScore, and human/LLM evaluations. Our qualitative analysis confirms QUIDS’ effectiveness in generating accurate intent descriptions for exploratory search. Our work contributes to improving the interaction between users and search engines by providing feedback to the user in exploratory search settings. [^1] Our code is available at https://github.com/menauwy/QUIDS
QUIDS: Query Intent Description for Exploratory Search
via Dual Space Modeling
Yumeng Wang Leiden University Leiden, The Netherlands y.wang@liacs.leidenuniv.nl          Xiuying Chen MBZUAI Abu Dhabi, United Arab Emirates xiuying.chen@mbzuai.ac.ae          Suzan Verberne Leiden University Leiden, The Netherlands s.verberne@liacs.leidenuniv.nl
1 Introduction
@@
Figure 1: A user-facing application of query intent generation in exploratory search. The system’s inferred intent is generated by contrasting top-ranked (pseudo-relevant) and low-ranked (pseudo-irrelevant) documents. Key information contributing to the inferred intent is shown in bold.
In exploratory search (Palagi et al., 2017), users often issue vague or underspecified queries to investigate unfamiliar topics through iterative refinement. This process gives rise to a persistent usability challenge, which we call the dual-blind problem: Users are uncertain about how to express their information needs; as a result they formulate ambiguous queries; the system silently infers the user’s intent based on these ambiguous queries, without providing explicit feedback and retrieves mixed-quality results at best. This leads to a self-reinforcing cycle of the user receiving results that are misaligned with their actual information need. This cycle is difficult to break with traditional query understanding methods Li et al. (2023); Lu et al. (2024); Wang et al. (2023) that operate in the pre-retrieval stage. Their goal is to optimize ranking effectiveness, not to provide feedback to the user. They offer little transparency about how the system arrived at its results, which is especially a problem when users are unsure of their intent.
In this work, we study the task of generating a natural-language query intent description (Zhang et al., 2020) that reflects what the system likely inferred the query to mean. The description is generated in the post-retrieval stage, incorporating system-inferred relevance of documents. These descriptions are not mechanistic explanations of the ranker, but instead serve as user-facing proxies of the system’s inferred intent. By contrasting top-ranked (pseudo-relevant) and low-ranked documents, the intent descriptions provide feedback that helps users identify mismatches between their intended and inferred query meanings. This feedback supports more effective query refinement and improves the overall search experience.
Figure 1 illustrates an application scenario of query intent generation in exploratory search. Since the ground truth intent behind the query is unknown to the search system, it relies on the retrieved documents to infer the query intent. From the top-ranked documents (considered relevant by the search engine), key terms like ‘Freon-12’, ‘refrigerant’, and ‘phased out’ are captured and emphasized in the intent description. In contrast, topics such as ‘costs’ and ‘prices’, which appear in both high-ranked and low-ranked documents, are excluded from the final intent description. The resulting description provides diagnostic feedback on how the system understood the query, helping users assess whether the retrieved results align with their latent intent.
We introduced a novel dual-space modeling approach, QUIDS, for the query intent generation task. It models query intent through dual-space contrastive learning by performing contrastive learning in two complementary spaces, explicitly separating intent-relevant and irrelevant semantic information. Specifically, the method consists of: (i) a representation space via dual encoders and (ii) a novel disentangling space in the decoder. This dual-space design enables the model to subtract irrelevant semantics from relevant ones, generating concise and more accurate intent descriptions. Furthermore, we propose an intent-driven hard negative sampling strategy to expand the irrelevant representation space and improve contrastive learning during training.
Experiments on the Q2ID benchmark (Zhang et al., 2020), including TREC and SemEval datasets, show that our model significantly outperforms strong baselines, including the prior Q2ID-specific method, LLM-based, and Query-focused Summarization methods, both in automatic and human evaluations. Qualitative analysis confirms that QUIDS effectively filters out distracting or misleading content and generates concise intent descriptions. Our contributions are: (i) Our model generates high-quality intent descriptions, with performance significantly enhanced by incorporating hard negative data augmentation during training. (ii) We introduce contrastive learning in both the representation space and the disentangling space of transformer models, effectively capturing contrasting information from relevant and irrelevant documents. (iii) We perform a thorough evaluation of our model, providing us with insights into the model’s strengths and weaknesses, as well as its potential application scenarios, especially for exploratory search.
2 Related Work
2.1 Query Understanding
Our work is related to traditional query understanding tasks such as classification Broder (2002); Verberne et al. (2013), clustering Wen et al. (2002); Hong et al. (2016), and expansion Wang et al. (2023); Mo et al. (2023); Jagerman et al. (2023). However, unlike these methods, which operate in the pre-retrieval stage to optimize retrieval effectiveness, Zhang et al. (2020) proposes the Query-to-Intent-Description (Q2ID) task in the post-retrieval stage that aims to generate search systems’ inferred intent of a user query based on both relevant and irrelevant documents. Unlike their method, we directly model a query-aware irrelevant intent space via dual-space contrastive learning, and enhance the performance with hard negative data augmentation, leading to a more precise intent description.
2.2 Query-focused Summarization
In settings where annotated intent descriptions are available, a related task to Q2ID is query-focused summarization (QFS) (Vig et al., 2022; Pagnoni et al., 2023). QFS is a subtask of text summarization that aims to generate a summary of one or multiple documents, guided by a query. Traditional methods rely on unsupervised extraction, ranking text segments by similarity and query relevance (Wan and Xiao, 2009; Feigenblat et al., 2017). Recent QFS datasets (Kulkarni et al., 2020; Fabbri et al., 2022; Zhong et al., 2021) have enabled the rise of QA-driven approaches (Su et al., 2020, 2021). More advanced techniques model query relevance through evidence ranking (Xu and Lapata, 2021), latent query optimization (Xu and Lapata, 2022), or pipeline architectures like the coarse-to-fine model in Xu and Lapata (2020). To handle long documents, extract-then-abstract strategies (Vig et al., 2022) use sparse attention and segment scoring. Other innovations include question-driven pretraining (Pagnoni et al., 2023), contrastive learning (Sotudeh and Goharian, 2023), and joint token-utterance modeling with query-aware attention (Liu et al., 2023a).
We use QFS models as baselines for the Q2ID task, but there is a fundamental difference between QFS and Q2ID: QFS aims to compress the content of retrieved documents to help users consume information, whereas Q2ID aims to generate a description of what the system likely inferred about the query intent, based on retrieval results. We provide a comparison table with related tasks in Appendix A.
3 Methods
@@
Figure 2: Overview of our proposed pipeline. From left to right, we show (a) Intent-Driven Negative Augmentation method, (b) Contrastive decoder structure with dual cross-attention layers, (c) and (d) Contrastive learning via dual space modeling.
3.1 Pipeline Framework
We define the contrastive intent generation task as follows. Given a dataset $\mathcal{D}=\{(q,\mathcal{R},\mathcal{I},y)_{j}\}$ with $L$ samples, where $j\in\{0,1,...,L\}$: $q$ is a query, $\mathcal{R}=\{r_{1},r_{2},...,r_{|\mathcal{R}|}\}$ is a collection of relevant documents for the query, $\mathcal{I}=\{i_{1},i_{2},...,i_{|\mathcal{I}|}\}$ is a collection of irrelevant documents and $y$ is the human-annotated ground truth query intent. The modeling goal is to learn the distinctions between relevant and irrelevant inputs based on a query, while generating a system-inferred intent description that exclusively highlights the relevant aspects related to the query. To achieve this, our training pipeline consists of 2 steps: (1) Intent-Driven Negative Augmentation (IDNA) and (2) Dual Space Modeling (DualSM).
3.2 Intent-Driven Negative Augmentation (IDNA)
The purpose of IDNA is to mine hard negative documents as irrelevant documents from the entire dataset $D$ based on the query, its relevant document collections, and the ground truth intent, i.e.,
$IDNA(q,\mathcal{R},y,\mathcal{D})=\mathcal{I}^{\prime}$
where $\mathcal{I}^{\prime}=\{i^{\prime}_{1},i^{\prime}_{2},...,i^{\prime}_{h}\}$ with $h$ the expected number of irrelevant documents. Inspired by Liu et al. (2022) on choosing in-context sample strategies for in-context learning, we design a method to choose intent-aware hard negative samples based on semantic similarity. Specifically, we use a Sentence Transformer model (Reimers and Gurevych, 2019) to represent both positive and negative samples from the training data in a vector space. Then we choose negative samples close to the positive ones in this space. The negative sample augmentation makes the task more challenging for the model, hence improving its discriminative capabilities. As shown in Figure 2 (a), we augment hard negative samples for each training query using Algorithm 1.
Algorithm 1 Intent-Driven Negative Augmentation (IDNA)
1:query $q$, relevant document collection $\mathcal{R}$, irrelevant document collection $\mathcal{I}$, whole dataset document corpus $\mathcal{D}$, target size $S$, threshold $\tau$
2:augmented irrelevant documents $\mathcal{I}^{\prime}$ for $q$
@@ 3.2 Intent-Driven Negative Augmentation (IDNA)
[…]
10:while $|\mathcal{I}^{\prime}|<S$ do
11:  Sample a document $d$ from $\mathcal{D}$
12:  if $\cos(h_{R^{*}},\text{Encode}(d))>\tau$ then
13:   Add $d$ to $\mathcal{I}^{\prime}$
14:  end if
15:end while
For encoding, we use a Sentence Transformers model pre-trained on the MSMARCO Passage Ranking dataset (Nguyen et al., 2016). In practice, we set 0.8 to the similarity threshold with an analysis of its effect in subsection 5.4.3.
3.3 Dual Space Modeling (DualSM)
@@
Dual Space Modeling aims to contrastively generate a descriptive intent for the query by modeling query-aware relevant and irrelevant intent spaces:
$DualSM(q,\mathcal{R},i^{\prime})=\hat{y}$
@@
We use a Transformer-based encoder-decoder architecture, with the BART-large model (Lewis et al., 2020) and the T5-large model Raffel et al. (2020a) as backbones, as illustrated in Figure 2(b). To capture the relationship between a query and its relevant and irrelevant documents, we implement a Siamese dual encoder architecture. Based on the encoder outputs, contrastive learning is performed in the representation space to differentiate between embeddings for relevant and irrelevant documents. Correspondingly, we design a contrastive decoder to model query-aware relevant and irrelevant intent spaces in a disentangled manner.
3.3.1 Representation Space Modeling (RSM)
We design a dual cross-encoder architecture to distinguish relevant and irrelevant documents for a given query by jointly encoding each query-document pair. Relevant documents, which often share similar topics tied to the query’s intent, are concatenated based on their ranking (Section 3.2, Step 1) and encoded together. In contrast, irrelevant documents can be irrelevant to a query in diverse ways, making it impractical to model a meaningful and comprehensive irrelevant feature representation space. Therefore, we focus on a single irrelevant document $i^{\prime}$ at each training step, using a hard negative sample from the augmented irrelevant document collection $\mathcal{I^{\prime}}$.
To model the feature space, we project the encoder’s final hidden states through a linear layer. Document embeddings are obtained via average pooling over token representations. We optimize the representation space by pulling relevant embeddings closer and pushing the irrelevant one away (Figure 2(c)). The objective is to minimize:
$\mathcal{L}_{rel}=\textstyle\sum^{k}_{m=1}\sum^{k}_{n=m+1}d(e_{m},e_{n})$ (1)